# Unveiling the Hidden: Movie Genre and User Bias in Spoiler Detection

### v2

@@ Abstract.
Spoilers in movie reviews are important on platforms like IMDb and Rotten Tomatoes, offering benefits and drawbacks. They can guide some viewers’ choices but also affect those who prefer no plot details in advance, making effective spoiler detection essential. Existing spoiler detection methods mainly analyze review text, often overlooking the impact of movie genres and user bias, limiting their effectiveness. To address this, we analyze movie review data, finding genre-specific variations in spoiler rates and identifying that certain users are more likely to post spoilers. Based on these findings, we introduce a new spoiler detection framework called GUSD  [^1] The code is available at https://github.com/AI-explorer-123/GUSD (Genre-aware and User-specific Spoiler Detection), which incorporates genre-specific data and user behavior bias. User bias is calculated through dynamic graph modeling of review history. Additionally, the R2GFormer module combines RetGAT (Retentive Graph Attention Network) for graph information and GenreFormer for genre-specific aggregation. The GMoE (Genre-Aware Mixture of Experts) model further assigns reviews to specialized experts based on genre. Extensive testing on benchmark datasets shows that GUSD achieves state-of-the-art results. This approach advances spoiler detection by addressing genre and user-specific patterns, enhancing user experience on movie review platforms.
@@
Spoiler Detection, Movie Genre, User Bias, Mixture-of-Experts
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Spoilers in movie reviews have become an important component of the movie-viewing experience on popular platforms like IMDb and Rotten Tomatoes (Cao et al., 2019). For those who hope to learn the plot of the movie in advance to judge whether they like it or not, the spoilers are helping, while for those who prefer to experience a movie without prior knowledge, the spoilers can severely diminish the enjoyment by revealing crucial plot points, undermining suspense, and eliciting negative emotions among viewers (Loewenstein, 1994). Thus, effective spoiler detection methods are crucial for maintaining a positive user experience.
@@
Figure 1. An illustrative example of the data used in our spoiler detection study. The image shows a review of the movie Moana, including the movie’s genres (Adventure, Animation), summary, the review’s content, and user-specific details. All reviews from the user are color-coded: blue indicates non-spoiler content, while red indicates spoiler content.
@@
Existing spoiler detection methods primarily focus on the textual content of reviews. For example, DNSD (Chang et al., 2018) integrates review sentences and movie genres, while SpoilerNet (Wan et al., 2019) utilizes a Hierarchical Attention Network and incorporates the item-specificity information. More recent approaches, such as MVSD (Wang et al., 2023), incorporate advanced techniques like syntax-aware graph neural networks and external movie knowledge to improve detection performance. Nevertheless, these methods still exhibit notable limitations. Solely relying on textual content proves insufficient for robust spoiler detection (Wang et al., 2023). Moreover, spoilers are often genre-specific, with varying characteristics depending on the movie’s genre — for instance, suspense films focus on plot details, whereas action films emphasize fight scenes. As a result, two significant challenges in spoiler detection remain unaddressed:
@@
Figure 2. (a) Spoiler rate across different movie genres (partial) in LCS dataset. (b) Kernel density estimation plot and distribution histogram of spoiler across different users.
@@
Figure 3. Overview of our proposed GUSD framework, which integrates genre-specific information, user behavior bias, and global perceptive RetGAT for spoiler detection. It preprocesses movie, user, and review data with MLP and language models, and captures user bias via dynamic graph modeling. Then the data is processed by the R2GFormer component. An Aggregator merges these features, and then GMoE assigns reviews to experts based on genres. Finally, a classifier detects spoilers using the aggregated features.
2. Related Work
2.1. Spoiler Detection
@@
The goal of automatic spoiler detection is to identify spoilers in reviews from domains like television (Boyd-Graber et al., 2013), books (Wan et al., 2019), and movies (Boyd-Graber et al., 2013). Existing approaches to spoiler detection can be broadly classified into three categories: keyword matching methods, machine learning techniques, and deep learning models.
2.1.1. Keyword matching methods.
@@
These approaches rely on a set of predefined keywords to identify spoilers. Examples include keywords related to sports teams or events (Nakamura and Tanaka, 2007), or actors’ names (Golbeck, 2012). Although useful in specific scenarios, this method requires manual keyword definition and lacks generalizability across different application contexts.
2.1.2. Machine learning techniques.
@@
These methods often involve topic modeling or support vector machines using handcrafted features. For example, Guo et al. (Guo and Ramakrishnan, 2010) applied a bag-of-words representation combined with an LDA-based model for spoiler detection. Jeon et al. (Jeon et al., 2013) developed an SVM classifier incorporating four extracted features, while Boyd et al. (Boyd-Graber et al., 2013) utilized lexical features and meta-data of review subjects (e.g., movies and books) in an SVM model.
2.1.3. Deep learning models.
@@
These models mainly leverage NLP techniques, employing RNNs, LSTMs, Transformer, and language models to process review texts and movie information through end-to-end training. Bao et al. (Bao et al., 2021) utilized LSTMs, BERT, and RoBERTa for sentence-level spoiler detection. DNSD (Chang et al., 2018) focused on incorporating external genre information using GRU and CNN. SpoilerNet (Wan et al., 2019) introduced item-specificity and bias with bi-RNN enhanced by GRU. SDGNN (Chang et al., 2021) leveraged dependency relations between context words in sentences with graph neural networks to capture semantics.
@@
While some existing methods incorporate genre information and user bias  (Chang et al., 2018; Wróblewska et al., 2021), they often rely on quite simple techniques such as concatenating or adding these additional features to the initial review features. Such approaches lack the sophistication needed for more effective and intricate modeling of genre features and user biases.
2.2. Mixture of Experts
@@
The Mixture of Experts (MoE) approach, grounded in the Divide-and-Conquer principle, segments an input sample into sub-tasks and trains specialized experts for each sub-task. This method is extensively utilized in NLP to boost model capacity (Shazeer et al., 2017) and enhance reasoning capabilities (Madaan et al., 2021). Shazeer et al. (Shazeer et al., 2017) introduced a sparsely-gated Mixture-of-Experts layer, enabling conditional computing in large language models. Fedus et al. (Fedus et al., 2022) developed simplified routing algorithms for MoE to enhance training stability and reduce computational costs. Furthermore, Soft-MoE (Puigcerver et al., 2023) was introduced to mitigate issues like training instability and token dropping inherent in traditional MoE approaches.
@@
3. Methodology
@@
3.1. Data preprocessing
3.1.1. Meta information.
For review nodes, user nodes, and movie nodes, each type possesses meta information. After collecting the meta information for all three types of nodes, we pad them to the same length. A two-layer MLP is then employed as the meta encoder, producing the meta embeddings $N_{m}$.
3.1.2. Textual information.
@@ 3.1. Data preprocessing
3.1.3. User bias acquisition using dynamic-graph pre-training.
To better capture the dynamic information attributes of users, we adopt the dynamic graph to handle users’ review history flexibly. Thus, we need to convert the static dataset into a dynamic format. We utilize the given connections between different nodes and the time information of the reviews to form the dynamic event stream. The details about the formation of the graph will be displayed in Section 3.2. Considering the absence of additional information about the dynamic edges, we simply initial the edge features as zero vectors. Then, we employ the robust DyGFormer (Yu et al., 2023) as our dynamic graph encoder to capture the interactions among various nodes and utilize Link Prediction as the downstream task. We specifically obtain the features of user nodes as user bias $U_{b}$ from the dynamic graph encoder.
3.2. R2GFormer
@@
3.2.2. RetGAT.
Inspired by the work of (Sun et al., 2023; Fan et al., 2024; Nikolentzos et al., 2020), we propose RetGAT, which extends the RetNet framework by integrating a global perception into GAT, incorporating explicit exponential decay for nodes within $k$ hops and truncation for nodes beyond $k$ hops. This method ensures a broad receptive field while balancing performance and computational complexity by dynamically adjusting the influence of nodes based on their distance, with closer nodes having a higher impact on the final node features.
To achieve this global receptive field, we utilize $k$ parallel GAT layers to separately aggregate information from $k$-hop neighbors. For each layer, node features are aggregated from different $k$-hop neighborhoods, applying decay factors to control the influence of information from various hops. Beyond $k$ hops, the influence of nodes is truncated to maintain computational efficiency and focus on relevant information within the $k$-hop range. Note that we correct the algorithm that previous work (Atwood and Towsley, 2016; Wang et al., 2020) used to compute $k$-hop neighbors, and the details are available in the supplementary material.
@@
(1) $\delta_{h}$ $=\exp(-\alpha\cdot h),$
@@
(3) ${N_{g}^{(l)}}$ $=\text{AGGREGATOR}_{r}^{(l)}(\text{${N_{g,1}^{(l)}}$},\text{${N_{g,2}^{(l)}}$},\ldots,\text{${N_{g,h}^{(l)}}$},\ldots,\text{${N_{g,k}^{(l)}}$}).$
@@
(4) $g_{j}^{(l)}=\text{AGGREGATOR}^{(l)}_{g}(\{n_{g,i}^{(l)}\mid i\in\mathcal{N}_{j}\}),$
@@
(5) $[g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]=\text{TRM}([g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]),$
@@
(6) $z_{i}^{(l)}$ $=\text{MEAN}(\{g_{j}^{(l)}\mid j\in G_{i}\}),$
(7) $n_{g,i}^{(l)}$ $=\text{\text{MLP}}([n_{g,i}^{(l)}\parallel z_{i}^{(l)}]).$
@@
(8) $N_{g}^{(l)}=\text{R2GFormer}(N_{g}^{(l-1)},\mathcal{E},G).$
@@ 3.3. Multimodal Fusion
(9) $U_{g},U_{m},U_{b}=\text{TRM}([U_{g}\ U_{m}\ U_{b}]$ $,U=\text{\text{MLP}}([U_{g}\parallel U_{m}\parallel U_{b}])),$
(10) $R_{g},R_{m}=\text{TRM}([R_{g}\ R_{m}]$ $,R=\text{\text{MLP}}([R_{g}\parallel R_{m}])),$
(11) $M_{g},M_{m}=\text{TRM}([M_{g}\ M_{m}]$ $,M=\text{\text{MLP}}([M_{g}\parallel M_{m}]),$
@@
(12) $r_{i}=[\ r_{i}\parallel m_{i}\parallel u_{i}\ ].$
3.4. GMoE
@@
(13) $r_{i}=\text{AGGREGATOR}_{m}(\{\text{Expert}_{j}(r_{i})\mid\forall{j}\in{G_{i}}\}).$
@@ 3.5. Learning and Optimization
(14) $Loss=-\sum_{i\in\mathcal{R}}{y_{i}\log\hat{y_{i}}}+\lambda\sum_{\theta\in\Theta}{\theta^{2}}.$
@@
Table 1. Accuracy, AUC, and binary F1-score of GUSD and three types of baseline methods on two spoiler detection datasets. We run all experiments five times to ensure a consistent evaluation and report the average performance as well as standard deviation in parentheses. Bold indicates the best performance, underline the second best. GUSD consistently outperforms the three types of methods on both benchmarks.

### v3

@@ Abstract
Spoilers in movie reviews are important on platforms like IMDb and Rotten Tomatoes, offering benefits and drawbacks. They can guide some viewers’ choices but also affect those who prefer no plot details in advance, making effective spoiler detection essential. Existing spoiler detection methods mainly analyze review text, often overlooking the impact of movie genres and user bias, limiting their effectiveness. To address this, we analyze movie review data, finding genre-specific variations in spoiler rates and identifying that certain users are more likely to post spoilers. Based on these findings, we introduce a new spoiler detection framework called GUSD (Genre-aware and User-specific Spoiler Detection), which incorporates genre-specific data and user behavior bias. User bias is calculated through dynamic graph modeling of review history. Additionally, the R2GFormer module combines RetGAT (Retentive Graph Attention Network) for graph information and GenreFormer for genre-specific aggregation. The GMoE (Genre-Aware Mixture of Experts) model further assigns reviews to specialized experts based on genre. Extensive testing on benchmark datasets shows that GUSD achieves state-of-the-art results. This approach advances spoiler detection by addressing genre and user-specific patterns, enhancing user experience on movie review platforms. Our source code is available at https://github.com/AI-explorer-123/GUSD
Keywords:
@@
Spoiler Detection Movie Genre User Bias Mixture-of-Experts
1 Introduction
@@
Spoilers in movie reviews have become an important component of the movie-viewing experience on popular platforms like IMDb and Rotten Tomatoes [5]. For those who hope to learn the plot of the movie in advance to judge whether they like it or not, the spoilers are helping, while for those who prefer to experience a movie without prior knowledge, the spoilers can severely diminish the enjoyment by revealing crucial plot points, undermining suspense, and eliciting negative emotions among viewers [21]. Thus, effective spoiler detection methods are crucial for maintaining a positive user experience.
@@
Figure 1: An illustrative example of the data used in our spoiler detection study. The image shows a review of the movie Moana, including the movie’s genres (Adventure, Animation), summary, the review’s content, and user-specific details. All reviews from the user are color-coded: blue indicates non-spoiler content, while red indicates spoiler content.
@@
Existing spoiler detection methods primarily focus on the textual content of reviews. For example, DNSD [6] integrates review sentences and movie genres, while SpoilerNet [36] utilizes a Hierarchical Attention Network and incorporates the item-specificity information. More recent approaches, such as MVSD [38], incorporate advanced techniques like syntax-aware graph neural networks and external movie knowledge to improve detection performance. Nevertheless, these methods still exhibit notable limitations. Solely relying on textual content proves insufficient for robust spoiler detection [38]. Moreover, spoilers are often genre-specific, with varying characteristics depending on the movie’s genre — for instance, suspense films focus on plot details, whereas action films emphasize fight scenes. As a result, two significant challenges in spoiler detection remain unaddressed:
@@
Figure 2: (a) Spoiler rate across different movie genres (partial) in LCS dataset. (b) Kernel density estimation plot and distribution histogram of spoiler across different users.
@@
2 Related Work
2.1 Spoiler Detection
@@
The goal of automatic spoiler detection is to identify spoilers in reviews from domains like television [3], books [36], and movies [3]. Existing approaches to spoiler detection can be broadly classified into three categories: keyword matching methods, machine learning techniques, and deep learning models.
2.1.1 Keyword matching methods.
@@
These approaches rely on a set of predefined keywords to identify spoilers. Examples include keywords related to sports teams or events [25], or actors’ names [12]. Although useful in specific scenarios, this method requires manual keyword definition and lacks generalizability across different application contexts.
2.1.2 Machine learning techniques.
@@
These methods often involve topic modeling or support vector machines using handcrafted features. For example, Guo et al. [13] applied a bag-of-words representation combined with an LDA-based model for spoiler detection. Jeon et al. [16] developed an SVM classifier incorporating four extracted features, while Boyd et al. [3] utilized lexical features and meta-data of review subjects (e.g., movies and books) in an SVM model.
2.1.3 Deep learning models.
@@
These models mainly leverage NLP techniques, employing RNNs, LSTMs, Transformer, and language models to process review texts and movie information through end-to-end training. Bao et al. [2] utilized LSTMs, BERT, and RoBERTa for sentence-level spoiler detection. DNSD [6] focused on incorporating external genre information using GRU and CNN. SpoilerNet [36] introduced item-specificity and bias with bi-RNN enhanced by GRU. SDGNN [7] leveraged dependency relations between context words in sentences with graph neural networks to capture semantics.
@@
While some existing methods incorporate genre information and user bias  [6, 40], they often rely on quite simple techniques such as concatenating or adding these additional features to the initial review features. Such approaches lack the sophistication needed for more effective and intricate modeling of genre features and user biases.
2.2 Mixture of Experts
@@
The Mixture of Experts (MoE) approach, grounded in the Divide-and-Conquer principle, segments an input sample into sub-tasks and trains specialized experts for each sub-task. This method is extensively utilized in NLP to boost model capacity [33] and enhance reasoning capabilities [23]. Shazeer et al. [33] introduced a sparsely-gated Mixture-of-Experts layer, enabling conditional computing in large language models. Fedus et al. [10] developed simplified routing algorithms for MoE to enhance training stability and reduce computational costs. Furthermore, Soft-MoE [29] was introduced to mitigate issues like training instability and token dropping inherent in traditional MoE approaches.
@@
Figure 3: Overview of our proposed GUSD framework, which integrates genre-specific information, user behavior bias, and global perceptive RetGAT for spoiler detection. It preprocesses movie, user, and review data with MLP and language models, and captures user bias via dynamic graph modeling. Then the data is processed by the R2GFormer component. An Aggregator merges these features, and then GMoE assigns reviews to experts based on genres. Finally, a classifier detects spoilers using the aggregated features.
3 Methodology
@@
3.1 Data preprocessing
3.1.1 Meta information.
For review nodes, user nodes, and movie nodes, each type possesses metadata (details can be found in the supplementary material). After collecting the metadata for all three types of nodes, we pad them to the same length. A two-layer MLP is then employed as the meta encoder, producing the meta embeddings $N_{m}$.
3.1.2 Textual information.
@@ 3.1 Data preprocessing
3.1.3 User bias acquisition using dynamic-graph pre-training.
To better capture the dynamic information attributes of users, we adopt the dynamic graph to handle users’ review history flexibly. Thus, we need to convert the static dataset into a dynamic format. We utilize the given connections between different nodes and the time information of the reviews to form the dynamic event stream. The details about the formation of the graph will be displayed in Section 3.2. Considering the absence of additional information about the dynamic edges, we simply initialize the edge features as zero vectors. Then, we employ the robust DyGFormer [42] as our dynamic graph encoder to capture the interactions among various nodes and utilize Link Prediction as the downstream task. We specifically obtain the features of user nodes as user bias $U_{b}$ from the dynamic graph encoder.
3.2 R2GFormer
@@
3.2.2 RetGAT.
Inspired by the work of [34, 9, 26], we propose RetGAT, which extends the RetNet framework by integrating a global perception into GAT, incorporating explicit exponential decay for nodes within $k$ hops and truncation for nodes beyond $k$ hops. This method ensures a broad receptive field while balancing performance and computational complexity by dynamically adjusting the influence of nodes based on their distance, with closer nodes having a higher impact on the final node features.
To achieve this global receptive field, we utilize $k$ parallel GAT layers to separately aggregate information from $k$-hop neighbors. For each layer, node features are aggregated from different $k$-hop neighborhoods, applying decay factors to control the influence of information from various hops. Beyond $k$ hops, the influence of nodes is truncated to maintain computational efficiency and focus on relevant information within the $k$-hop range. Note that we correct the algorithm that previous work [1, 37] used to compute $k$-hop neighbors, and the details are available in the supplementary material.
@@
$\delta_{h}$ $=\exp(-\alpha\cdot h),$ (1)
@@
${N_{g}^{(l)}}$ $=\text{AGGREGATOR}_{r}^{(l)}(\text{${N_{g,1}^{(l)}}$},\text{${N_{g,2}^{(l)}}$},\ldots,\text{${N_{g,h}^{(l)}}$},\ldots,\text{${N_{g,k}^{(l)}}$}).$ (3)
@@
$g_{j}^{(l)}=\text{AGGREGATOR}^{(l)}_{g}(\{n_{g,i}^{(l)}\mid i\in\mathcal{N}_{j}\}),$ (4)
@@
$[g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]=\text{TRM}([g_{1}^{(l)}g_{2}^{(l)}\cdots g_{j}^{(l)}\cdots g_{c}^{(l)}]),$ (5)
@@
$z_{i}^{(l)}$ $=\text{MEAN}(\{g_{j}^{(l)}\mid j\in G_{i}\}),$ (6)
$n_{g,i}^{(l)}$ $=\text{\text{MLP}}([n_{g,i}^{(l)}\parallel z_{i}^{(l)}]).$ (7)
@@
$N_{g}^{(l)}=\text{R2GFormer}(N_{g}^{(l-1)},\mathcal{E},G).$ (8)
@@ 3.3 Multimodal Fusion
$U_{g},U_{m},U_{b}=\text{TRM}([U_{g}\ U_{m}\ U_{b}]$ $,U=\text{\text{MLP}}([U_{g}\parallel U_{m}\parallel U_{b}])),$ (9)
$R_{g},R_{m}=\text{TRM}([R_{g}\ R_{m}]$ $,R=\text{\text{MLP}}([R_{g}\parallel R_{m}])),$ (10)
$M_{g},M_{m}=\text{TRM}([M_{g}\ M_{m}]$ $,M=\text{\text{MLP}}([M_{g}\parallel M_{m}]),$ (11)
@@
$r_{i}=[\ r_{i}\parallel m_{i}\parallel u_{i}\ ].$ (12)
3.4 GMoE
@@
$r_{i}=\text{AGGREGATOR}_{m}(\{\text{Expert}_{j}(r_{i})\mid\forall{j}\in{G_{i}}\}).$ (13)
@@ 3.5 Learning and Optimization
$Loss=-\sum_{i\in\mathcal{R}}{y_{i}\log\hat{y_{i}}}+\lambda\sum_{\theta\in\Theta}{\theta^{2}}.$ (14)
@@
4 Experiment
4.1 Experiment Settings
4.1.1 Dataset.
To evaluate our GUSD framework along with 14 other representative baselines on two widely recognized datasets: LCS [38] and Kaggle [24]:
  • Kaggle, introduced in 2019, consists of 573,913 valid reviews, 263,407 users, and 1,572 movies. And about 25.87% (150,924) of the reviews are spoilers.
Note that both datasets include the genre information of all movies, with specific categories defined according to IMDb standards, which facilitates the operation of our GUSD framework. Following MVSD [38], we randomly split the reviews into training, validation, and test sets with a ratio of 7:2:1.
4.1.2 Baselines.
To achieve a comprehensive evaluation, we compare GUSD with pre-trained language models, GNN-based models, and task-specific baselines. For the pre-trained language models, the procedure involves feeding the review text into the model, averaging all token embeddings, and then applying two fully connected layers to perform spoiler detection. Regarding the GNN-based models, the graph neural network takes the output of RoBERTa [19] as the initial node features. Below, we provide a concise overview of each baseline method.
  • BERT [8] is a language model pre-trained on extensive natural language data, using masked language modeling and next sentence prediction tasks.
  • RoBERTa [19] improves upon BERT by eliminating the next sentence prediction task and enhancing masking techniques.
  • BART [18] is a pre-trained language model that advances traditional autoregressive models through bidirectional encoding and denoising objectives.
  • DeBERTa [14] refines BERT by implementing disentangled attention and an improved mask decoder, making it a more advanced language model.
  • Bge-Large [41] is trained on a comprehensive training dataset C-MTP, combining vast unlabeled data and diverse labeled data.
  • GCN [17] is a foundational graph neural network that performs convolutions on graph nodes and their neighbors, effectively propagating information.
  • R-GCN [32] extends GCN to handle multi-relational graphs by incorporating relation-specific weights.
  • GAT [35] is a graph neural network that applies attention mechanisms to dynamically assign importance to neighboring nodes.
  • SimpleHGN [22] is tailored for heterogeneous graphs, integrating multiple types of nodes and edges with a shared embedding space and adaptive aggregation strategies.
  • GPS [30] propose a recipe to build a general, powerful, scalable graph Transformer with linear complexity.
  • HGT [15] design node- and edge-type dependent parameters to characterize the heterogeneous attention over each edge for modeling Web-scale heterogeneous graphs.
  • DNSD [6] is a spoiler detection method that employs a CNN-based genre-aware attention mechanism.
  • SpoilerNet [36] uses a hierarchical attention network and GRU alongside item and user bias terms for spoiler detection.
  • MVSD [38] leverages external movie knowledge and user networks to detect spoilers.
Table 1: Accuracy, AUC, and binary F1-score of GUSD and three types of baseline methods on two spoiler detection datasets. We run all experiments five times to ensure a consistent evaluation and report the average performance as well as standard deviation in parentheses. Bold indicates the best performance, underline the second best. GUSD consistently outperforms the three types of methods on both benchmarks.